OpenStudy (anonymous):

Rational Functions: x-2/4 + x/3 = 3

OpenStudy (anonymous):

 First find LCM of 4 and 3

OpenStudy (anonymous):

Yes that's it

OpenStudy (anonymous):

So, what is the LCM of 4 and 3

OpenStudy (anonymous):

It's 1

OpenStudy (anonymous):

12

OpenStudy (anonymous):

yes it is 12

Parth (parthkohli):

I love when all the terms have the same denominators so I could equate the numerators. For example,\[\frac{2x}{5} + \frac{5}{4} = x \iff \frac{8x}{20} + \frac{25}{20} = \frac{20x}{20} \iff 8x + 25 = 20x\]

OpenStudy (anonymous):

what?

Parth (parthkohli):

It's just that you make ALL the terms you see in the equation with denominator 12.

OpenStudy (anonymous):

Could you write that?

Parth (parthkohli):

Well, I'd just give a part of it:\[\frac{3\cdot(x - 2)}{3 \cdot 4} + \frac{4\cdot x}{4\cdot 3} = \frac{3 \cdot 12}{1 \cdot 12}\]

Parth (parthkohli):

(BTW, that dot stands for multiplication)

OpenStudy (anonymous):

so the answer is 3

OpenStudy (anonymous):

36/12 = 3

Parth (parthkohli):

Dude, no, you gotta make all the numbers have a denominator of 12, then equate the numerators.

Parth (parthkohli):

Therefore,\[3(x - 12)+4x = 36\]